OpenStudy (anonymous):

if the equation a^2t^2 - act + a =0 has x and y as roots then (x-y)^2=?

OpenStudy (apoorvk):

If you notice, this is a quadratic in 'a'. that means, sum of roots: x+y = -ct/t^2 and product of roots = xy = 1/t^2 now, \[(x-y)^2 = x^2 + y^2 - 2xy = x^2 + y^2 + 2xy - 4xy = (x+y)^2 - 4xy\]

OpenStudy (anonymous):

dum of roots x + y = (ac) / a^2 = c/a xy = a / a^2 = 1 / a plug these values into what apoorvk got and you have your answer

OpenStudy (anonymous):

* sum of roots

OpenStudy (anonymous):

dumb of the roots sounds better

OpenStudy (anonymous):

lol

OpenStudy (apoorvk):

Lol!

OpenStudy (anonymous):

thxxxx

OpenStudy (anonymous):

yw